[figure: 9.7 cm]

MAGNETIZATION REVERSAL IN Fe/Si/Fe FILM WITH FERROMAGNETIC INTERLAYER EXCHANGE

A.B. Chizhik^{a}, S.L. Gnatchenko^{a}, M. Baran^{b}, K. Fronc^{b}, R. Szymczak^{b} and

R. Zuberek^{b}

aInstitute for Low Temp. Phys. and Engin., NAS of Ukraine, 61103 Kharkov, Ukraine bInstitute of Physics, PAS, Al. Lotnikow 32/46, 02-668 Warsaw, Poland

Process of magnetization reversal was studied in Fe/Si/Fe film at 300 K and 10 K. The Fe30A/Si14A/Fe30A sandwich was prepared by dc sputtering. The experiments were performed using the magneto-optical Kerr effect technique and SQUID magnetometer. The ferromagnetic resonance measurements have shown the presence of ferromagnetic and biquadratic interlayer exchange as well as cubic and in-plane uniaxial anisotropy. The directions of easy axes of in-plane uniaxial anisotropy differ in iron layers. The jumps of magnetization in the magnetic field, which are accompanied by the local hysteresis takes place in the magnetization hysteresis loops. Different numbers of jumps were observed depending on the direction of the magnetic field in the film plane and on the temperature. The experimental results were analyzed based on the model taking into account the ferromagnetic and biqudratic exchange as well as cubic and uniaxial anisotropy. Numerical calculations of the field dependencies of the magnetization have been performed for room and helium temperature. It was found that the observed jumps of magnetization are associated with the overcoming of the energetic barriers related to hard axes in the two iron layers. The spontaneous and field induced magnetic configurations are determined as a result of comparison of experiment and calculation.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : A. B. Chizhik

Address for correspondence : Institute for Low Temperature Physics and Engineering Academy of Sciences of Ukraine 61103 Kharkov Ukraine

Email address : chizhik@ilt.kharkov.ua